# An Empirical Study of Developer-Provided Context for AI Coding Assistants in Open-Source Projects

### v1

@@ 4. Context Types in Cursor Rules
Code Style Conventions for the code style, including naming, formatting, code writing, and other stylistic choices. Favor the use of functional components over class components. 0.65    0.05
@@
QA Practices for managing, preventing, and reporting errors, including logging, exception handling, and test requirements for verification. Comprehensive testing: Test all integration features. 0.68    0.10
@@
Behavior Directs LLMs to follow specific rules and guidelines when generating responses, e.g., to reference other documentation, use tools, or add verification steps. Always ask clarifying questions when you don’t have full context or understanding of a task or question. 0.36    0.02
[…]
Env. Specifies the technology stack (e.g., libraries, tools, hardware) and commands used to build, initialize, run, or test the project. Run the CLI ‘init‘ command: ‘npx trigger.dev@latest init‘. 0.72    0.07
@@
From the qualitative analysis, we identified five high-level categories of context types that developers commonly provide: Project, Convention, Guideline, LLM Directives, and Examples. The taxonomy of context types is illustrated in Figure 2. Detailed definitions, representative examples of each context type, and their usage distributions with explanations are provided in Section 4. In the following sections, we describe each category and its subcategories in detail, with representative examples.
@@ 4.1. Project
Environment. Many cursor rules describe the technology stack, including the programming languages, frameworks, libraries, and tools used in the project. They also outline the setup and configuration steps required to run the project, such as environment variables, configuration files, and necessary initialization commands. Most of this information overlaps with content typically found in the installation section of software documentation. The main purpose appears to be constraining the LLM to generate code compatible with the existing technology stack. Detailed explanations of the setup and commands can help LLM agents understand how to run tests, build the project, or deploy it.
@@ 4.2. Convention
The convention category captures rules that prescribe specific conventions or standards to follow when contributing to a software project. It is usually tied to a project-selected style guide or specific project usage requirement.
@@
Language/Framework Specific. Language/framework usage includes conventions related to the specific programming languages, frameworks, or libraries used in the project, such as preferred idioms, patterns, or best practices. For example, “Use WordPress ‘@wordpress/element‘ instead of direct React import.”
@@ 4.3. Guideline
Quality Assurance. The most frequently mentioned code in the guideline was quality assurance. These rules provided recommendations on testing, code reviews, debugging, and other quality assurance practices to ensure the reliability and maintainability of the software. For example, “Utilize guard clauses to handle preconditions and invalid states early.”
General Programming. Many cursor rules include very high-level programming best practices, such as separation of concerns, modularity, type safety, and code reuse, which any experienced developer would already be familiar with. For example, “Maintain module separation of concerns”.
@@
4.4. LLM Directives
Another category we detected is LLM Directives, which captures context that provides specific instructions and directives for LLMs to follow when generating code snippets. Unlike the previous three categories, which contain more general information that could be helpful for any developer working on the project and have also been widely discussed in prior literature on software documentation or developer communication practices (maalej2013patterns; freund2015contextualizing), this category specifically targets the capabilities of LLMs and provides guidance tailored to them.
@@ 4.4. LLM Directives
Behavior. As the most frequently mentioned code in the LLM Directives category, the behavior code provides instructions on how the LLM should behave when generating code snippets, such as being cautious, avoiding assumptions, verifying information, and asking clarifying questions. In more details, developers often uses a variety of strategies to guide LLM behavior in their projects. One common approach is structured reasoning, where prompts explicitly direct the model’s thought process using techniques such as chain-of-thought prompting (wei2023chain; kojima2023large) or self-consistency (wang2023self), helping the model arrive at more reliable outputs. Another frequent practice is asking for clarification: when prompts are ambiguous or underspecified, instructions may tell the LLM to request additional context, for instance, “Always ask clarifying questions when you don’t have full context or understanding of a task or question.” Many developers also include self-verification steps, requiring the LLM to check its assumptions, dependencies, or constraints before generating code, such as “Before generating any code, you MUST verify: 1. Are you importing from @trigger.dev/sdk/v3? If not, STOP and FIX.” To prevent errors or undesired outcomes, instructions sometimes include specific prohibitions or antipatterns. Finally, developers frequently embed project-specific rules that are non-generalizable but essential for their particular context, for example, “Do not run test-teardown even if tests are successful—I will do that manually or request it explicitly.” These practices collectively help ensure that LLMs behave predictably, safely, and in alignment with project requirements.
@@
Workflow. The next most commonly mentioned theme in the LLM Directives category was workflow, which specifies a structured, multistep process or sequence of tasks that an LLM should follow to complete a task, often by mimicking real-world human software development processes. Many of these workflows are designed to decompose complex tasks into smaller, manageable steps, such as “Before responding to any request, follow these steps: 1. Request Analysis, …, 2. Solution Planning, …”. Some rules are written as conditional sequences, under which certain steps are executed only when specific conditions are met, for example, “When a user request begins with CHANGELOG:, add a changelog entry based on the provided prompt in the following files:”.
@@
Persona. Many cursor rules specified the persona the LLM should adopt, such as “You are an expert in Python, writing an AI application called chatgpt-mirai-qq-bot. It is a workflow-based chatbot system.” Along with these persona definitions, some rules provide a high-level overview of the entire project, which the LLM is expected to be familiar with. We also observed that most persona instructions directed the LLM to assume technical roles, such as software engineer, with fewer projects emphasizing non-technical personas, such as personal assistants or friends. Comparing to behavior and workflow codes, persona instructions were less frequently mentioned in cursor rules.
@@
Output Formatting and Granularity are used to provide instructions on how the agent should respond in terms of format and level of detail. Formatting mainly directs the LLM on how to structure its responses, such as specifying templates, required formats, languages (e.g., “Respond in Chinese”), or inclusion of references. Granularity provides instructions on the level of detail the LLM should include in its responses, such as “Deliver the response in a minimal yet complete form. Avoid unnecessary verbosity and tangential remarks.”
4.5. Examples
@@
There is one auxiliary content category called Examples, which captures different types of examples provided, usually alongside established rules, such as naming, formatting, or usage examples. For example, “Example: https://github.com/brainlid/langchain/pull/261. The documentation literature (maalej2013patterns) has highlighted the importance of examples, especially the code examples, in helping developers understand and apply guidelines effectively. While we found the similar trend in cursor rules, the proportion of examples is represented relatively low in our quantitative results, because we programmatically excluded code blocks from the analysis (Section 3.3). Therefore, the number of example codes should be treated as a lower bound.
@@ 4.6. Discussion
Overall, our qualitative analysis revealed that developers provide a diverse range of context types in cursor rules, spanning project-specific information, coding conventions, high-level guidelines, LLM-specific instructions, and examples. The prevalence of project-related context highlights developers’ emphasis on conveying the unique aspects of their software projects to LLMs, which is crucial for generating relevant and accurate code snippets. The widespread use of conventions and guidelines indicates developers’ desire to maintain consistency and quality in code contributions, aligning with established best practices. The presence of LLM-specific instructions, although less common, underscores the growing recognition of the need to tailor LLM behavior to better suit the requirements of software development tasks.
We also found that the spectrum of context types, except for LLM-specific instructions, largely aligns with established best practices for software documentation and developer communication (maalej2013patterns; freund2015contextualizing). This suggests that developers are leveraging cursor rules as an extension of traditional documentation, adapting familiar practices to the context of LLM-assisted coding.
@@ 5. Characteristics of Cursor Rules
In Section 3.3, we identified different types of context that developers commonly provide in cursor rules. During this process, we observed that the context types may vary depending on various project characteristics, such as programming language and application domain. In this section, we further investigate these variations by exploring the following research questions, which are motivated by our observations during the coding process and by related literature on developer communication practices (maalej2013patterns; freund2015contextualizing), project characteristics (plinfluencecontext), and LLM prompt engineering strategies (sahoo2024systematic; schulhoff2024prompt).
@@
Prior research (plinfluencecontext; plinfluencecontext1) has shown that different programming languages influence developers’ communication practices and the types of context they prioritize. For instance, statically typed languages such as Java, Rust, and TypeScript may require less explicit context regarding type safety, as their type systems enforce correctness at compile time. In contrast, dynamically typed languages such as Python and JavaScript could require more detailed context related to type checking and error handling, since such issues are only detectable at runtime. Furthermore, for recently adopted programming languages such as Rust (rustneeddifferent; rustneeddifferent1), developers may need to provide additional context about language-specific features and idiomatic practices. In this question, we investigated whether these language-specific differences are actually reflected in the context provided in cursor rules.
@@
RQ3: To what extent do developers provide multiple types of context in cursor rules?
Many guidelines for traditional documentation recommend that developers provide multiple types of information to give a comprehensive overview of the project. For example, Maalej and Robillard (maalej2013patterns) suggest that effective documentation should include both high-level overviews and low-level implementation details to support various developer needs. Similarly, many studies on documentation have found that developers face challenges when documentation lacks code examples or explanations of the rationale behind design decisions (sohan2017study; ebert2021exploratory). In RQ3, we investigated whether developers follow these best practices when authoring cursor rules.
RQ4: To what extent do developers reuse cursor rules?
In software engineering, reusability is a key principle that promotes efficiency and consistency (copypaste). Developers are encouraged to reuse existing code snippets, libraries, and frameworks to expedite development and maintain consistency across projects. Similarly, prior work on prompt engineering has highlighted the benefits of reusing effective prompts from external sources or previous projects to enhance LLM performance (prompt11). In our analysis, we observed that developers often reuse context from existing documentation, coding standards, and best practices when authoring cursor rules. Additionally, we found several community-maintained repositories providing template cursor rules, which encourage the reuse of existing cursor rules. With this research question, we aimed to understand the extent to which developers reuse cursor rules from external sources or prior projects when authoring their own rules.
@@
Since writing documentation for LLMs is a relatively new practice, we hypothesized that the timing of when developers add cursor rules to their projects may influence the type of context they provide. For example, if a repository was created long before cursor rules were introduced, developers might have reuse project-related information by copying from existing documentation in the repository, rather than writing detailed and diverse context specifically for LLMs. Conversely, for newly created repositories, developers could provide more high-level and LLM-specific instructions to guide LLMs in generating code from scratch. Thus, we aimed to understand how the timing of cursor rule addition relative to repository creation affects the context provided.
To answer these research questions, we collected additional metadata for 401 repositories we analyzed in Section 3.3, and present our analysis and findings below.
@@ 5.1. RQ1: Programming Language
Findings. Figure 4 shows the distribution of the average number of codes for each context type by programming language. Overall, different programming languages tend to have different distributions of context types.
Developers tend to provide less context when working with statically typed languages such as Go, C#, and Java, and provide more context when working with dynamic languages like JavaScript and PHP. This suggests that developers expect that the stricter syntax and type checking in statically typed languages allow LLMs to infer more information directly from the code, reducing the need for additional context (typeisgood1; typeisgood). For example, repositories of JavaScript as main languages tend to provide more guideline-related context (Mean=92) and more conventions (M=34) while those using TypeScript, a statically typed superset of JavaScript, has a lower average number of guideline-related context (M=50) and conventions (M=34).
@@
Some less dominant and more domain specific programming languages, such as C#, tend to have more LLM-related context (M=31) provided in cursor rules. This may be because C# is often used in enterprise applications and game development, where LLMs need to be more cautious about the code they generate to avoid potential issues. Additionally, its similarity in structure to Java and other object-oriented languages—which are more commonly used and prevalent in LLM training data—may increase the necessity of providing additional context.
Among all programming languages, JavaScript and TypeScript are the most common languages for which developers provide examples (M=24 and M=21, respectively). This may be because JavaScript and TypeScript are often used for frontend development with a wide variety of frameworks and libraries that are updated at a fast pace, so developers feel the need to provide more examples to help LLMs understand specific usage patterns and best practices for these languages.
Limitations. Our analysis only considered the dominant programming language of each repository, as the GitHub API only provides a single primary language per repository. This may oversimplify the analysis for multi-language projects, such as web development projects that are full-stack and contain multiple programming languages. However, we believe the relative comparisons between programming languages still hold, as we applied the same criteria across all repositories.
Additionally, our dataset is skewed, with over 50% of repositories being Typescript or JavaScript. While this represents the current distribution of repositories adopting cursor rules on GitHub, more studies should be conducted to better understand the differences and requirements when working with LLMs across various programming languages.
@@ 5.2. RQ2: Application Domain
To identify application domains, we used the list of GitHub topics set by the repository owners via the official GitHub API. Considering that each repository can have multiple topics and can be assigned by owners freely, we first performed clustering over these topics. We removed generic terms such as “ai” and “open source,” and excluded non-English topics. Then, we used repositories that had at least two topics assigned (317 repositories) for further analysis. Then, we performed topic modeling with BERTopic (berttopic) on concatenated topics. We then used UMAP (umap) with the number of neighbors set to 15 for dimensionality reduction and HDBSCAN (hdbscan) for clustering. We chose HDBSCAN because it does not require pre-defining the number of clusters, and can identify noise points that do not belong to any cluster. We set the minimum cluster size as 20, 5% of the dataset, and a minimum sample size of 2 to enforce at least two repositories per topic. In addition, we performed a greedy search over different hyperparameter combinations, optimizing and balancing the number of topics, the meaningfulness of topics, the ratio of unclustered topics, and the coherence of individual items with topics (measured using the c_v metric with the Gensim model (c-v1; c-v2)), along with topic diversity (measured by the number of unique tokens over the entire document length), to determine the most appropriate clustering outcomes. The final selected number of topics was six. The selected combination of hyperparameter allowed us to achieve a low unclustered topic ratio (0.3%), acceptable topic coherence (0.52), and good topic diversity (0.93). The research team manually checked the meaningfulness of the topics by reviewing the top keywords and representative repositories for each topic. Since the c_v metric is highly dependent on the corpus (lim2024aligning), and the topics assigned by developers were usually quite diverse, we believe these metrics indicate a good clustering result. For the embedding model, we used the officially recommended “all-MiniLM-L6-v2” (all-MiniLM-L6-v2). We excluded any unclustered repositories from further analysis.
We identified six distinct topics from 317 repositories: : Web Development(70 repositories), representing general web projects. : LLM Agent(62 repositories), for projects focused on building or evaluating large language model agents. : Frontend Development(61 repositories), for projects using frontend languages or frameworks such as React or WordPress, including domain-specific areas like e-commerce. For example, Angular (angular/angular). : Data Analysis(50 repositories), for projects centered on performing or managing data analysis tasks. : UI Development(44 repositories), for projects focused on design work, design utilities, and user interfaces. For example, ant-design-web3 (ant-design/ant-design-web3). : MCP Tools(29 repositories), for projects related to plugin development for coding assistants.
@@
Front-end and visually focused implementation projects, such as data analysis, web development, and frontend development, tend to provide more examples for LLMs (M=43, M=13, and M=13, respectively). Since these projects are usually written in JavaScript or TypeScript, this aligns with our finding (Section 5.1) that JavaScript/TypeScript projects include more examples.
For domains that do not have many publicly available, up-to-date API, such as MCP code server projects; developers tend to provide more project context to help LLMs understand the requirements and constraints of the project (M=69) (apihard). In contrast, for domains with abundant online resources, such as UI design projects, developers tend to provide less project context (M=45).
Developers tend to provide more guideline-related context when working on frontend development for projects that require more complex and unpredictable tasks that demand careful consideration before execution, such as LLM agent projects and data analysis projects (M=64 and M=66, respectively).
@@
In addition, there is a strong correlation between application domain and programming language, as certain domains tend to use specific programming languages. For example, frontend development projects predominantly use JavaScript and TypeScript, while data analysis projects often utilize Python and R. This correlation may confound our analysis of context types by application domain, as the programming language itself also influences the context types provided (Section 5.1). However, this is common in software engineering research, and disentangling the effects of programming language and application domain is inherently challenging. We found new insights that were not directly explained by programming language alone, such as the higher amount of guideline-related context in LLM agent projects and data analysis projects, showing that application domain can also play a significant role in shaping context provision.
5.3. RQ3: Context Co-occurrence
To test whether developers tend to provide extensive context covering multiple aspects of the project when authoring cursor rules, we analyzed the co-occurrence of different context categories within individual repositories.
In addition, we also counted the number of unique codes in each repository to understand the coverage of different context types.
Figure 6. Venn diagram of context type co-occurrences in cursor rules.
Findings. Figure 6 shows the co-occurrence of different context category in cursor rules. In general, developers were motivated to provide comprehensive context for LLMs, with 149 repositories (37.16%) providing four actual content context categories in their cursor rules and 99 repositories (24.69%) providing all five context categories (including examples).
Developers prefer to write cursor rules in a style similar to traditional project or collaboration documentation, which typically covers best practices and standards for software development (prana2019categorizing; fronchetti2023contributing). In our case, this includes project, conventions, and guidelines category. People usually pay the same amount of attention to the tuple combinations of those three context types, which were relatively evenly distributed (Project and Convention: 296 repositories, 73.8%; Project and Guideline: 310 repositories, 77.3%; Guideline and Convention: 311 repositories, 77.6%). Examples are usually provided by the traditional documentation to illustrate other context types and help people use the other context types more effectively as well (exampleeffectiveness). In our case, the Example category is usually co-occurs with all other three categories: Guideline, LLM Directive, and Convention (172 repositories, 42.9%). Individually, examples are most frequently provided alongside Guidelines (193, 48.1%), Conventions (189, 47.1%), LLM Directives (187 repositories, 46.6%)
Some developers choose to only focus on specific aspects when writing cursor rules to let LLM focus on a really specific aspects while generating code. 32 repositories (7.98%) provided only one category of context in their cursor rules; most of them were for guideline-related context (11 repositories, 2.74%).

### v2

@@ 4. Context Types in Cursor Rules
[…]
Change Describes shifts or updates in a project’s technical implementation or configuration. No need for ‘postcss-import‘ or ‘autoprefixer‘ anymore 0.17    0.01
@@
Structure Conventions for how files and directories are created, named, and organized within a software project. Use the ‘apps‘ directory for Next.js and Expo applications. 0.59    0.04
@@
UI Principles to improve usability, accessibility, and internationalization (i18n). All interactive elements (links, buttons, form controls, custom components) must be operable via a keyboard. 0.34    0.02
@@
Granularity Directives instructing an LLM to adjust response detail. Avoid unnecessary verbosity and tangential remarks. 0.12    0.00
@@
From our qualitative analysis, we identified five high-level context types commonly provided by developers: Project, Convention, Guideline, LLM Directive, and Example. Detailed definitions, representative examples of each context type, and their usage distributions are shown in Section 4. In the following sections, we describe each category and its subcategories in detail, with representative examples.
@@ 4.1. Project
Environment. Many cursor rules describe the technology stack, including the programming languages, frameworks, libraries, and tools used. They also outline configuration steps, such as environment variables, configuration files, and required initialization commands. Most of this information overlaps with content typically found in the setup section of software documentation. The main purpose appears to be constraining the LLM to generate code compatible with the existing technology stack and helping LLM agents understand how to run tests, build the project, or deploy it.
@@ 4.2. Convention
The convention category includes rules that prescribe standards or practices for contributing to a project, typically based on a project-selected style guide or specific project usage requirement.
Language/Framework Specific. Language/framework usage includes conventions related to the specific programming languages, frameworks, or libraries used in the project, such as preferred idioms, patterns, or best practices. For example, “Use WordPress ‘@wordpress/element’ instead of direct React import.”
@@ 4.3. Guideline
Quality Assurance. The most frequently mentioned code in the guideline was quality assurance, which includes instructions on testing, code reviews, debugging, and other practices to ensure the reliability and maintainability of the software. For example, “Utilize guard clauses to handle preconditions and invalid states early.”
General Programming. Many cursor rules include very high-level programming best practices, such as separation of concerns, modularity, and type safety, which any experienced developer already knows. For example, “Maintain module separation of concerns”.
@@
Performance is used to provide guidance on efficient resource usage, such as “Cache expensive operations.” Consistency instructs agents to align with existing patterns, styles, and structures, such as “Follow existing patterns…” UI focuses on usability, accessibility, and internationalization, such as “All interactive elements must be operable via a keyboard.” Security addresses practices to mitigate vulnerabilities and protect data integrity, such as “Implement proper nonce verification”. And dependency provides guidance in package usage to ensure maintainability, such as “Ensure packages are properly isolated and dependencies are correctly managed.”
4.4. LLM Directive
Another category we identified is LLM Directives, which captures context that provides instructions and directives for LLMs to follow when responding. Unlike the previous three categories, which provide general information useful to any developer and have been widely discussed in prior literature (maalej2013patterns; freund2015contextualizing), this category specifically targets LLM capabilities and offers tailored guidance.
@@ 4.4. LLM Directive
Behavior. As the most frequently mentioned code in the LLM Directives category, the behavior code provides instructions on how the LLM should behave when responding, such as being cautious, avoiding assumptions, verifying information, and asking clarifying questions. In more details, developers often uses a variety of strategies to guide LLM behavior in their projects. One common approach is structured reasoning, where prompts explicitly direct the model’s thought process using techniques such as chain-of-thought prompting (wei2023chain; kojima2023large) or self-consistency (wang2023self), helping the model arrive at more reliable outputs. Another frequent practice is asking for clarification: when prompts are ambiguous or underspecified, instructions may tell the LLM to request additional context, for instance, “Always ask clarifying questions when you don’t have full context or understanding of a task or question.” Many developers also include self-verification steps, requiring the LLM to check its assumptions, dependencies, or constraints before generating code, such as “Before generating any code, you MUST verify: 1. Are you importing from @trigger.dev/sdk/v3? If not, STOP and FIX.” To prevent errors or undesired outcomes, instructions sometimes include specific prohibitions or antipatterns. Finally, developers frequently embed project-specific rules that are non-generalizable but essential for their particular context, for example, “Do not run test-teardown even if tests are successful—I will do that manually or request it explicitly.” These practices collectively help ensure that LLMs behave predictably, safely, and in alignment with project requirements.
@@
Workflow. The next most commonly mentioned theme was workflow, which specifies a structured, multistep process or sequence of tasks that an LLM should follow to complete a task, often by mimicking real-world human software development processes. Many of these workflows are designed to decompose complex tasks into smaller, manageable steps, such as “Before responding to any request, follow these steps: 1. Request Analysis, …, 2. Solution Planning, …”. Some rules are written as conditional sequences, under which certain steps are executed only when specific conditions are met, for example, “When a user request begins with CHANGELOG:, add a changelog entry based on the provided prompt in the following files:”.
@@
Persona. Many cursor rules specify the persona the LLM should adopt, such as “You are an expert in Python, writing an AI application called chatgpt-mirai-qq-bot. It is a workflow-based chatbot system.” Along with these persona definitions, some rules provide a high-level overview of the entire project, which the LLM is expected to be familiar with. We also observed that most persona instructions direct the LLM to assume technical roles, such as software engineer, with fewer projects emphasizing non-technical personas, such as story maker. Compared to behavior and workflow codes, persona instructions are less frequently mentioned in cursor rules.
@@
Output Formatting and Granularity are used to instruct the agent on how to respond in terms of format and level of detail. Formatting primarily directs the LLM on how to structure its responses, such as specifying templates, required formats, languages (e.g., “Respond in Chinese”), or the inclusion of references. Granularity provides instructions on the level of detail the LLM should use in its responses, such as “Deliver the response in a minimal yet complete form. Avoid unnecessary verbosity and tangential remarks.”
4.5. Example
@@
We also identified one auxiliary content category, Example, which are typically provided alongside established rules, such as naming, formatting, or usage examples. For instance, “Example: https://github.com/brainlid/langchain/pull/261. maalej2013patterns highlighted the importance of examples in helping developers. While we found a similar trend in cursor rules, the proportion of examples is relatively low in our quantitative results because code blocks were excluded (Section 3.3). Therefore, the number of example codes should be considered a lower bound.
@@ 4.6. Discussion
Overall, our qualitative analysis reveals that developers provide a diverse range of context types in cursor rules, spanning project-specific information, coding conventions, high-level guidelines, LLM-specific instructions, and examples. The prevalence of project-related context underscores developers’ focus on conveying their software’s unique aspects to LLMs. The widespread use of conventions and guidelines indicates developers’ desire to maintain consistency and quality, aligning with established best practices. The presence of LLM-specific instructions, although less common, highlights the growing need to tailor LLM behavior to better suit the requirements of software development tasks.
We found that the spectrum of context types, aside from LLM-specific instructions, largely aligns with established software documentation practices (maalej2013patterns; freund2015contextualizing). This suggests that developers are using cursor rules as an extension of traditional documentation, adapting familiar practices to the context of LLM-assisted coding.
@@ 5. Characteristics of Cursor Rules
In Section 4, we identified different types of context that developers commonly provide in cursor rules. During this process, we observed that the context types may vary depending on various project characteristics, such as programming language and application domain. In this section, we further investigate these variations by exploring the following research questions, which are motivated by our observations during the coding process and by related literature on developer communication practices (maalej2013patterns; freund2015contextualizing), project characteristics (plinfluencecontext), and LLM prompt engineering strategies (sahoo2024systematic; schulhoff2024prompt).
@@
Prior research (plinfluencecontext; plinfluencecontext1) has shown that different programming languages influence developers’ communication practices and the types of context they prioritize. For instance, statically typed languages such as Java, Rust, and TypeScript may require less explicit context regarding type safety, as their type systems enforce correctness at compile time. In contrast, dynamically typed languages such as Python and JavaScript could require more detailed context related to type checking and error handling, since such issues are only detectable at runtime. Furthermore, for recently adopted programming languages such as Rust (rustneeddifferent; rustneeddifferent1), developers may need to provide additional context about language-specific features and idiomatic practices. In this question, we investigated whether these language-specific differences are actually reflected in those rules.
@@
RQ3: To what extent do developers reuse cursor rules?
In software engineering, reusability is a key principle that promotes efficiency and consistency (copypaste). Developers are encouraged to reuse existing code snippets, libraries, and frameworks to expedite development and maintain consistency across projects. Similarly, prior work on prompt engineering has highlighted the benefits of reusing effective prompts from external sources or previous projects to enhance LLM performance (prompt11). In our analysis, we observed that developers often reuse rules from existing documentation, coding standards, and best practices when authoring cursor rules. Additionally, we found several community-maintained repositories providing template cursor rules, which encourage the reuse of existing cursor rules. With this research question, we aimed to understand the extent to which developers reuse cursor rules from external sources or prior projects when authoring their own rules.
RQ4: To what extent do developers provide multiple types of context in cursor rules?
Many guidelines for traditional documentation recommend that developers provide multiple types of information to give a comprehensive overview of the project. For example, Maalej and Robillard (maalej2013patterns) suggest that effective documentation should include both high-level overviews and low-level implementation details to support various developer needs. Similarly, many studies on documentation have found that developers face challenges when documentation lacks code examples or explanations of the rationale behind design decisions (sohan2017study; ebert2021exploratory). In RQ4, we investigated whether developers follow these best practices when authoring cursor rules.
@@
Since writing documentation for LLMs is a new practice, we hypothesized that the timing of cursor rule addition may influence the type of context included. In older repositories, developers might reuse existing project documentation rather than create detailed context for LLMs, whereas in newer repositories, they may provide more high-level, LLM-specific instructions to guide code generation. Therefore, we want to understand how the timing of rule addition relative to repository creation affects the context provided.
@@ 5.1. RQ1: Programming Language
Findings. Different programming languages exhibit different distributions of context types, as shown in fig. 5.
Developers tend to provide less context when working with statically typed languages such as Go, C#, and Java, and provide more context when working with dynamic languages like JavaScript and PHP. This suggests developers expect stricter type checking in statically typed languages helps LLMs infer more from the code, reducing the need for extra context (typeisgood1; typeisgood). For example, repositories of JavaScript as main languages tend to provide more guideline-related context (Mean=92) and more project-related context (M=102), while those using TypeScript, a statically typed superset of JavaScript, has a lower average number of guideline-related context (M=50), and project-related context (M=59).
@@
Some less dominant, domain specific languages, such as C#, tend to have more LLM-related context (M=31) in cursor rules. This may be because C# is often used in enterprise applications and game development, where LLMs need to be more cautious about the generated code to avoid potential issues. Additionally, its similarity in structure to Java and other object-oriented languages—which are more commonly used and prevalent in LLM training data—may increase the necessity of providing additional context.
Among all languages, JavaScript and TypeScript most commonly include examples in cursor rules (M=24 and M=21). This likely reflects their use in frontend development with rapidly evolving frameworks and libraries, leading developers to provide examples to help LLMs capture specific usage patterns and best practices.
Limitations. Our analysis considered only the dominant programming language of each repository, as the GitHub API provides a single primary language. This may oversimplify the analysis of multi-language projects, such as full-stack web development projects. Additionally, our dataset is skewed, with over 50% of repositories using TypeScript or JavaScript. Nevertheless, relative comparisons across languages should still hold, and this reflects the current adoption of cursor rules on GitHub. Further studies are necessary to better understand the differences and requirements when working with LLMs across programming languages.
@@ 5.2. RQ2: Application Domain
To identify application domains, we used the list of GitHub topics via the official GitHub API. Since each repository can have multiple topics and owners can assign them freely, we performed clustering over these topics using BERTopic (berttopic). Before clustering, we removed generic terms such as “ai” and “open source”, excluded non-English topics, and filtered out repositories labeled with only a single topic, resulting in 317 repositories for clustering.
To find the optimal clustering model, we conducted a greedy search over BERTopic hyperparameters to balance the number of topics, topic interpretability, the proportion of unclustered items, item-topic coherence (measured using the c_v metric with Gensim (c-v1; c-v2)), and topic diversity (measured as the ratio of unique tokens to total document length). The final model produced six topics, achieving a low unclustered ratio (0.3%), acceptable coherence (0.52), and high topic diversity (0.93). The research team further validated topic meaningfulness by manually inspecting top keywords and representative repositories. Given that c_v is corpus-dependent (lim2024aligning) and developer-assigned topics are highly diverse, these results indicate a satisfactory clustering outcome.
Finally, we used BERTopic with UMAP (umap) (15 neighbors) for dimensionality reduction and HDBSCAN (hdbscan) for clustering, which identifies noise points without requiring a predefined number of clusters. We set the minimum cluster size to 20 (5% of the dataset) and the minimum samples to 2 to ensure at least two repositories per topic. For embeddings, we used the officially recommended all-MiniLM-L6-v2 model (all-MiniLM-L6-v2). With excluding any unclustered repositories, we identified six distinct topics: : Web Development(70 repositories), general-purpose web projects. : LLM Agents(62 repositories), projects that build or evaluate LLM-based agents. : System Infrastructure(61 repositories), full-stack development and system design, including domain-specific applications such as e-commerce (e.g., OpenCut). : Data Analysis(50 repositories), projects focused on data analysis and data management tasks. : Frontend Development(44 repositories), frontend tools, design systems, and user interface development (e.g., ant-design-web3). : MCP Tools(29 repositories), plugins and extensions for coding assistants.
@@
More complex projects, such as data analysis, web development, and system infrastructure building, tend to provide more examples for LLMs (M=43, M=13, and M=13). These projects often involve intricate workflows, diverse data sources, complex component interactions, and extensive use of existing libraries, which may be difficult for LLMs to fully grasp without concrete examples.
For domains with few publicly available, up-to-date APIs (e.g., MCP code server projects), developers provide more project context to help LLMs understand project requirements (M = 69) (apihard), whereas for domains with abundant online resources (e.g., web development), developers provide less project context (M = 39).
Developers tend to provide more guideline-related context when working on projects that involve unpredictable or difficult-to-test tasks requiring careful consideration before execution, such as LLM agent projects and data analysis projects (M=64 and M=66).
@@
Application domain and programming language are inherently strongly correlated. For example, frontend projects predominantly use JavaScript or TypeScript, whereas data analysis projects often rely on Python or R. This correlation may confound our analysis of application domains, as programming language also influences the types of context provided (Section 5.1). However, such confounding is common in software engineering research, and disentangling the effects of programming language and application domain is inherently challenging. Despite this threat, we observed insights not solely attributable to programming language, such as the greater prevalence of guideline-related context in LLM agent and data analysis projects, showing the value of conducting a separate analysis on application domains.
5.3. RQ3: Proportion of Duplicated Lines
Findings. Overall, we identified a total of 19,917 (28.70%) duplicated lines across all repositories with an average of 49.67 duplicate lines (SD = 164.20) per repository. Figure 8 shows the percentage of code that are duplicated by context category.
Figure 5 shows the distribution of the average number of duplicated lines for each context type by programming language. We found that different programming languages exhibit different copying practices when writing cursor rules. For example, developers using programming languages for desktop and enterprise applications, such as C#, which require high domain specificity, tend to replicate cursor rules from similar projects, showing the highest average number of duplicate lines in project-related code (M=61.14, SD=161.77). In contrast, developers using languages for frontend applications, such as JavaScript, often copy pieces from various sources, resulting in evenly distributed duplication across all context types.
Figure 6. Percentage of code that are duplicated by context category.
Figure 7. Venn of context type co-occurrences.